OpenStudy (anonymous):

Find the specified value: S5=31/4 , r=1/2 , a1=? Please explain how to solve this.

OpenStudy (raden):

is it the GP ?

OpenStudy (raden):

if yes, use the formula : Sn = a(1 - r^n)/(1-r)

OpenStudy (raden):

so, S5 = 31/4 a(1 - r^n)/(1-r) = 31/4 a(1 - (1/2)^5)/(1- 1/2) = 31/4 a(1- 1/32)/(1/2) = 31/4 a(31/32) * 2 = 31/4 a (31/16) = 31/4 a = (31/4)/(31/16) a = ...